OpenStudy (anonymous):

The speed of light is 3x10*8 m/s. How long does it take light to reach the earth from the sun, which is 1.5x10^11 m away?

OpenStudy (anonymous):

time = distance/speed

OpenStudy (mayankdevnani):

always remember :- \[\huge time=\frac{distance}{speed}\] and time=? distance=1.5x10^11 m and speed=3x10*8 m/s. now plug these values in eq. we get, \[\huge time=\frac{1.5 \times10^{11}}{3 \times10^8 }\] now solve it

OpenStudy (mayankdevnani):

can you solve it? @danielle02

OpenStudy (anonymous):

500 s?

OpenStudy (anonymous):

@mayankdevnani is it 500 s ?

OpenStudy (mayankdevnani):

yes, you are right

OpenStudy (mayankdevnani):

@danielle02 good job